OpenStudy (anonymous):

If sin theta=3/5 and theta is in quadrant 2 determine in exact form sin theta+pi over 6

OpenStudy (loser66):

We have formula sin(a+b) = sin a *cos b + sin b *cos a your a = theta your b= pi/6 find cos theta, then plug all into it. 

OpenStudy (anonymous):

@Loser66 So I did all that and got the answer .9196 but I am off by a point when I check if it is right tho

OpenStudy (loser66):

show your work, please

OpenStudy (anonymous):

Alright the end work I got 3sqrt3 over 10 + 2/5 but it doesn't come out correct -___-

OpenStudy (loser66):

cos (theta) = 4/5 hence sin (theta + pi/6) = sin (theta) cos (pi/6) + sin (pi/6) cos (theta) = (3/5) (sqrt 3/2) + (1/2) (4/5) = 0.573

OpenStudy (anonymous):

But I still keep getting .919... answer ohmygod this is so frustarting lol @Loser66

OpenStudy (anonymous):

And when you go to check the answer it is suppose to be .90166

OpenStudy (freckles):

that is what it says when I punch in what loser has above

OpenStudy (freckles):

and 0.9196 is also what I'm getting

OpenStudy (freckles):

i think you are right @shortyyshayy

OpenStudy (freckles):

and they have an error

OpenStudy (anonymous):

@freckles Idk this is coming out weird cuz when checking it I am like a .1 off lol

OpenStudy (freckles):

just to be sure you are asked to evaluate \[\sin(\theta+\frac{\pi}{6})\]

OpenStudy (freckles):

well the other thing is it says it wants the EXACT form not an approximation

OpenStudy (freckles):

exact form would be \[\frac{2}{5}+\frac{3 \sqrt{3}}{10}\]

OpenStudy (freckles):

not .9196

OpenStudy (freckles):

which is just an approximation

OpenStudy (freckles):

omg I'm so dumb

OpenStudy (freckles):

I didn't see it said in quadrant 2

OpenStudy (freckles):

cos is negative there

OpenStudy (freckles):

\[\frac{3}{5} \frac{\sqrt{3}}{2}+\frac{-4}{5}\frac{1}{2}\]

OpenStudy (freckles):

which actually gives us .1196 so that still doesn't give us the answer you want

OpenStudy (loser66):

yes!!! @freckles

OpenStudy (anonymous):

where did you get -4/5 from tho??

OpenStudy (freckles):

it says theta is in quadrant 2

OpenStudy (freckles):

cos is negative there and sin is positive there

OpenStudy (freckles):

sin(theta)=3/5 cos(theta) can be 4/5 or -4/5 depending on where theta is

OpenStudy (freckles):

and since cos is negative then cos(theta)=-4/5

OpenStudy (freckles):

here is another way to look at it we are given sin(theta)=3/5 \[\cos^2(\theta)+\sin^2(\theta)=1 \\ \cos^2(\theta)+(\frac{3}{5})^2=1 \\ \cos^2(\theta)+\frac{9}{25}=1 \\ \cos^2(\theta)=1-\frac{9}{25} \\ \cos^2(\theta)=\frac{25}{25}-\frac{9}{25} \\ \cos^2(\theta)=\frac{16}{25} \\ \cos(\theta)=\pm \sqrt{\frac{16}{25}} \\ \cos(\theta)=\pm \frac{4}{5}\]

OpenStudy (freckles):

now deciding whether to use 4/5 or -4/5 entirely depends on where theta is

OpenStudy (freckles):